Hello!
in an election, 2.8 million votes were cast and each vote was for either Candidate 1 or Candidate 2. Candidate 2 received 28,000 more votes than Candidate 1. What percent of the 2.8 million votes were cast for Candidate 1?

thank you

Answer by jim_thompson5910(35256)

x = # of votes for candidate 1 (this number is in millions)
y = # of votes for candidate 2 (this number is in millions)


First equation:

x+y = 2.8

since a total of "2.8 million votes were cast and each vote was for either Candidate 1 or Candidate 2"

-------------------------------------------------------

28,000 = 28,000/1,000,000 = 0.028 million

Second equation:

y = x + 0.028

because "Candidate 2 received 28,000 more votes than Candidate 1"

this is the same as saying "Candidate 2 received 0.028 million more votes than Candidate 1"

-------------------------------------------------------

We'll use substitution to solve for x.


x+y = 2.8

x+x + 0.028 = 2.8 ... replace y with x + 0.028

2x+0.028 = 2.8

2x = 2.8-0.028

2x = 2.772

x = 2.772/2

x = 1.386


Remember that x = # of votes for candidate 1 (this number is in millions)

Since x = 1.386, this means that Candidate 1 got 1.386 million votes.

Divide this by 2.8 (the total): 1.386/2.8 = 0.495

Multiply by 100: 0.495*100 = 49.5%

So Candidate 1 got 49.5% of the vote.

Candidate 2 got 100 - 49.5 = 50.5% of the vote. This means Candidate 2 is the winner. This part can be easily determined from the fact that "Candidate 2 received 28,000 more votes than Candidate 1". Despite this, it helps to see the percentages as well.


Let me know if you need more help or if you need me to explain a step in more detail.
Feel free to email me at jim_thompson5910@hotmail.com
or you can visit my website here: http://www.freewebs.com/jimthompson5910/home.html

Thanks,

Jim